OpenStudy (anonymous):

I understand that infinity is a concept, however if one were to do: ∞+1, the answer would remain ∞. Would the same apply to doing ∞-1?

OpenStudy (amistre64):

purple + 3 = purple right?

OpenStudy (amistre64):

infinity is not a number ... so dont use it as a number it is a direction to travel towards

OpenStudy (anonymous):

Well you can't even say that \[\infty = \infty + 1 \] for reasons @amistre64 mentioned above. Technically neither of your statements are valid. What you may be saying is: \[\lim_{n \rightarrow \infty} n+1=\lim_{n \rightarrow \infty} n-1=\lim_{n \rightarrow \infty} n=\infty\] At which point you may notice that the concept of infinity is really dependent on your understanding of a limit.

OpenStudy (anonymous):

Very interesting, thanks. However, if we imagine a hotel with an infinite amount of rooms, and each room is filled with a guest, surely the hotel can accomodate an extra guest by asking each guest to move along one room? Thus still having an infinite amount of guests?

OpenStudy (anonymous):

I don't see how your analogy goes against what I said earlier.

OpenStudy (anonymous):

Derp, very true. Many thanks. Is there a way I can reward both of you with points, or can I only shoose one?

OpenStudy (amistre64):

i have enough medals; you can choose vf :)

OpenStudy (anonymous):

THERE CAN ONLY BE ONE.

OpenStudy (anonymous):

lol thx, amistre.

OpenStudy (anonymous):

Done, thanks guys. :)